OpenStudy (anonymous):

simplify: 3\sqrt{2}-\sqrt[3]{x}+7\sqrt{2}+6\sqrt[3]{x} anyone plzz

OpenStudy (saifoo.khan):

\[3\sqrt{2}-\sqrt[3]{x}+7\sqrt{2}+6\sqrt[3]{x} \] \[10 \sqrt 2 + 5\sqrt[3]{x}\]